OpenStudy (anonymous):

Find the value of x. A. 30 B. 60 C. 90 D. 120

OpenStudy (jjuden):

how can i find it with no graph or equation

OpenStudy (jjuden):

can u give me the problem so i can work it out please :)

OpenStudy (anonymous):

@jjuden

OpenStudy (anonymous):

@wendylisette

OpenStudy (anonymous):

hold up

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

but it isnt an answer choice

OpenStudy (solomonzelman):

Supplementary angles, So y-x+y+x=180 y=20 and 2x=y-x 2x=90-x x=90

OpenStudy (solomonzelman):

Made a mistake at first, sorry.

OpenStudy (anonymous):

its cool thanks you guy @SolomonZelman reaslly helped :) you to @wendylisette

OpenStudy (anonymous):

(:

OpenStudy (solomonzelman):

\(\bbox[5pt,cyan,border:2px solid blue]{•‿•}\)